A conducting ring of radius R is given a charge Q on its surface. Find the tension produced in the ring??

Grade:11

1 Answers

Vikas TU
6 years ago
When we accuse the wire of the given charge q=+Q. that will make pressure on it,that strain called circle strain. When we accuse the wire of the another charge q0=+Q 'which will be put at the focal point of the ring, the wire strain will be expanded to keep this portion into equillibrium once more. 
Consider a basic strip subtending calculated Ѳ at its inside. 
The string is in equillibrium in X-direction.For equillibrium in y-course, 
Fe = 2T sin dѳ/2… … (1) 
q/2π xdѳ 
Charge on the strip 
F = 1/4πɛ * q/2π x( dѳ*qo)/r2 - (2) 
Substituting 2 in 1 
2T dѳ/2 =1/4πɛ * q/2π x( dѳ*qo)/r2 
T =qq0/8π2ɛr2 N.